OpenStudy (anonymous):

can someone help me with square root property 5x^2-3=0 I need a solution set not sure how to do it

OpenStudy (anonymous):

5x^2 = 3 x^2 = 3/5 \[x = \pm \sqrt{\frac{3}{5}}\]

OpenStudy (anonymous):

So it is \[\sqrt{3}/5 and -\sqrt{3}/5\]

OpenStudy (anonymous):

if that square root sign includes the 5 as well as the three then yes

OpenStudy (anonymous):

that is what I had just making sure

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

no problem